# Statistical Inference for Temporal Difference Learning with Linear Function Approximation

### v1

@@ 3.3 Online estimation of the covariance matrix
Compared with the bootstrap-based procedure of Samsonov et al., (2024), our plug-in estimator $\widehat{\bm{\Lambda}}_{T}$ differ in three fundamental ways. (i) The upper bound in Theorem 3.4 vanishes, as a function of $T$, at the rate of $O(T^{-\frac{1}{3}})$, which is faster than the $O(T^{-\frac{1}{4}}\log T)$ convergence rate obtained from the bound (26). (ii) The leading term in the upper bound in Theorem 3.4 entails a smaller problem-related coefficient than that of (26). (iii) Perhaps most importantly, in order to estimate the bootstrap distribution $\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T}$ with an $O(T^{-\frac{1}{4}})$ error rate, it is necessary to store the sample trajectory $\mathcal{S}_{T}$ and generate $O(T^{\frac{1}{2}})$ perturbed trajectories. Consequently, the bootstrap procedure has a time complexity of order $O(T^{\frac{3}{2}}d^{2})$ and a space complexity of order $O(T)$. When $d\ll T$, this results in a significant computational border compared to our online procedure of Algorithm 1.
3.4 Statistical inference for $\bm{\theta}^{\star}$
By combining the results from Sections 3.2 and 3.3, we obtain a general methodology for constructing convex confidence sets for $\bm{\theta}^{\star}$ with guaranteed finite sample coverage, as shown in our final result, whose proof is in Appendix B.6.
@@ 3.4 Statistical inference for $\bm{\theta}^{\star}$
Consider TD with Polyak-Ruppert averaging (9) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, for all $t$ and with $\alpha=2/3$.For a given significance level $\delta\in(0,1)$, it holds that
@@
uniformly over all convex sets $\mathcal{C}_{\delta}\subset\mathbb{R}^{d}$ satisfying
$\mathbb{P}_{\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}_{d})}\left(\overline{\bm{\theta}}_{T}-\frac{\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}}{\sqrt{T}}\in\mathcal{C}_{\delta}\bigg{|}\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T}\right)\geq 1-\delta,$ (27)
where the variance matrix estimator $\widehat{\bm{\Lambda}}_{T}$ is given in (22).
It is important to highlight that the above coverage guarantee holds uniformly over any choice of the convex set $\mathcal{C}_{\delta}$ meeting condition (27). In particular, this class includes all hyper-rectangles (cartesian products of univariate intervals), returning simultaneous confidence intervals that are especially useful due to their simplicity and interpretability. The construction of $\mathcal{C}_{\delta}$ requires computing appropriate quantiles of the $\mathcal{N}(\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T})$ distribution, conditionally on $\overline{\bm{\theta}}_{T}$ and $\widehat{\bm{\Lambda}}_{T}$, depending on the choice of convex set. For the case of hyper-rectangles, this task is reduced to finding the quantile of the maximum of Gaussian variables.
Input: MDP transition kernel $\mathcal{P}$, stationary distribution $\mu$, discount factor $\gamma$, feature matrix $\bm{\Phi}$, initial stepsize $\eta_{0}$, stepsize decaying parameter $\alpha\in(0.5,1)$, sample size $T$
1 Initialize $\bm{\theta}=\bm{0}$, $\overline{\bm{\theta}}=\bm{0}$, $\overline{\bm{A}}=\bm{0}$, $\overline{\bm{AA}}=\bm{0}$, $\overline{\bm{Ab}}=\bm{0}$, $\overline{\bm{bb}}=\bm{0}$;
2 for $t=0:(T-1)$ do
3       sample $s\sim\mu$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s)$;
4       compute
$\bm{A}\leftarrow\bm{\phi}(s)[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})];$
$\bm{b}\leftarrow r(s)\bm{\phi}(s);$
5       perform TD update
$\bm{\theta}\leftarrow\bm{\theta}-\eta_{0}t^{-\alpha}(\bm{A}\bm{\theta}-\bm{b});$
6       perform online updates
$\overline{\bm{\theta}}\leftarrow\overline{\bm{\theta}}+\frac{1}{t}(\bm{\theta}-\overline{\bm{\theta}});$
$\overline{\bm{A}}\leftarrow\overline{\bm{A}}+\frac{1}{t}(\bm{A}-\overline{\bm{A}});$
$\overline{\bm{AA}}\leftarrow\overline{\bm{AA}}+\frac{1}{t}(\bm{A}\otimes\bm{A}-\overline{\bm{AA}});$
$\overline{\bm{Ab}}\leftarrow\overline{\bm{Ab}}+\frac{1}{t}(\bm{A}\otimes\bm{b}+\bm{b}\otimes\bm{A}-\overline{\bm{Ab}});$
$\overline{\bm{bb}}\leftarrow\overline{\bm{bb}}+\frac{1}{t}(\bm{b}\otimes\bm{b}-\overline{\bm{bb}});$
7compute
$\widehat{\bm{\Gamma}}\leftarrow\overline{\bm{AA}}(\overline{\bm{\theta}}\otimes\overline{\bm{\theta}})-2\overline{\bm{Ab}}(\overline{\bm{\theta}})+\overline{\bm{bb}};$
8 reshape
$\widehat{\bm{\Gamma}}\leftarrow\mathsf{reshape}(\widehat{\bm{\Gamma}},(d\times d))$
compute
Output: averaged TD estimator $\overline{\bm{\theta}}$, covariance matrix estimator $\widehat{\bm{\Lambda}}$
4 Numerical experiments
@@
We illustrate our theoretical results with some numerical experiments. In order to test our proposed framework for statistical inference, we consider a member of a family of MDPs that are designed to be difficult to distinguish among each other. The specifics of the MDP follow the precedent of Li et al., 2024b and we refer readers to Appendix D for the construction details. We examine the high-probability convergence of the norm of TD errors, the rate by which they converge to their asymptotic Gaussian distributions, the precision of our estimator for the covariance matrix, and the validity of our statistical inference procedure. While the full details and results of the experiments are postponed to Appendix D, we demonstrate some important findings in the simulations that verify our theoretical conclusions.
4.1 Uncertainty quantification with different choices of $\alpha$
In order to explore the role played by the parameter $\alpha$ that decides the decaying speed of the stepsizes, we run TD iterates using different choices of $\alpha$ on the MDP with feature dimension $d=3$, record the $95\%$ quantile of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ and measure the difference between the empirical and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$  [^3] For computational simplicity, this difference between distributions is not measured by the distance on convex sets as we used in Theorems 3.2 and 3.3. Details are showin in Appendix D. .
Figure 2(a) shows the decay of the $95\%$ quantile of the $L_{2}$ norm of $\overline{\bm{\Delta}}_{T}$; the slope of this log-log plot verifies our theoretical claim in Theorem 3.1 that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of $O(T^{-\frac{1}{2}})$ with high probability. Notably, the convergence rate is similar for different choices of $\alpha>\frac{1}{2}$ when $T$ is large. Figure 2 (b) reveals that the empirical distribution of TD estimation error converges to the asymptotic Gaussian distribution slowly when $\alpha$ is set to ${1}/{2}$ as suggested by Samsonov et al., (2024). Both plots appear to indicate that choosing $\alpha=\frac{1}{2}$ yields a slower converging rate both in $\ell_{2}$ norm and in distribution.
Refer to caption Refer to caption
@@ 4.1 Uncertainty quantification with different choices of $\alpha$
Figure 2: (a) $95\%$ quantile of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$, and (b) distance between empirical and asymptotic distributions of the TD estimation error $\overline{\bm{\Delta}}_{T}$, for an MDP with $d=3$ with different choices of $\alpha$.
4.2 Confidence intervals
Furthermore, we validate our statistical inference framework on the MDP with $d=3$ and TD learning with $\alpha=\frac{2}{3}$. We run 1000 independent trials of Algorithm 1, and at each iteration $T$, we construct both the individual and the simultaneous confidence intervals. Specifically, the individual confidence intervals for the three entries of $\bm{\theta}^{\star}$ are constructed by
@@ 5 Discussion
We develop high-probability convergence gaurantees and Berry-Esseen bounds for TD learning with linear function approximation and $i.i.d.$ samples in a non-asymptotic manner. We also propose an efficient way to construct a covariance estimator, thus paving the way for statistical inference on policy evaluation. One may naturally be interested in whether and how our techniques can be generalized to the more realistic scenario of Markov samples, where the TD estimation error is also gauranteed to be asymptotically normal (Fort, (2015)). On this front, the answer is multifaceted. For the high-probability convergence guarantee, our induction-based technique used in the proof of Theorem 3.1 may be generalized to Markovian samples under appropriate assumptions on the mixing time of the Markov chain. As for the CLT, a cornerstone result invoked in the proof of Theorem 3.2 holds only with $i.i.d.$ samples, and a new framework is necessary for generalizing the Berry-Esseen bound to Markovian samples. Furthermore, it would be of interest to derive specialized Berry-Esseen bounds for the sub-class of hyper-rectangles. In this setting, we expect to find a better dependence on the dimension $d$ of the feature space, possibly of poly-logarithmic, as opposed to polynmial, order. Finally, the plug-in estimator $\widehat{\bm{\Lambda}}_{T}$ may be generalized to a batch-mean estimator. Its efficiency, however, needs to be further scrutinized, as the variance estimation problem for samples generated from a Markov chain remains still largely unsolved.
Acknowledgements
W. Wu and A.Rinaldo are supported in part by NIH under Grant R01 NS121913. Gen Li is supported in part by the Chinese University of Hong Kong Direct Grant for Research. Y. Wei is supported in part by the NSF grants DMS-2147546/2015447, CAREER award DMS-2143215, CCF-2106778, CCF-2418156 and the Google Research Scholar Award.
Appendix A Preliminary facts
@@
A.1 Useful algebriac facts
The following lemma captures some basic properties in linear algebra regarding the relaitionship between the Frobenius and operator norms of matrices.
@@ A.1 Useful algebriac facts
    For any sequence of matrices $\bm{X}_{1},\bm{X}_{2},\ldots,\bm{X}_{n}\in\mathbb{R}^{d}$:
@@
The following lemma, introduced as Lemma 4.1 in Chen et al., (2020), bounds the difference between inverse matrices and would be useful for our analysis of the plug-in variance estimator.
@@
Let $\bm{X},\bm{Y}$ be two invertible matrices satifying $\|\bm{X}^{-1}(\bm{Y}-\bm{X})\|\leq\frac{1}{2}$. Then the difference between $\bm{Y}^{-1}$ and $\bm{X}^{-1}$ can be bounded by
@@
The following lemma reveals several critical algebriac features of the matrix $\bm{A}$.
@@
Let $\bm{A}_{t},\bm{A}$ be defined as in (9b) and (6a) respectively, and $\bm{\Sigma}$, $\lambda_{0}$ and $\lambda_{\Sigma}$ be defined as in (7) and (8). Then the following features hold:
@@
$\mathbb{E}[\bm{A}_{t}^{\top}\bm{A}_{t}]\leq\bm{A}+\bm{A}^{\top};$ (32)
$\left\|\bm{I}-\eta\bm{A}\right\|\leq 1-\frac{1-\gamma}{2}\lambda_{0}\eta,\quad\forall\eta\in\left(0,\frac{1+\gamma}{2\lambda_{\Sigma}}\right).$ (33)
@@
$-\left(\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top}\right)$
$\preceq\bm{\phi}(s)\bm{\phi}(s^{\prime})^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s)^{\top}$
$\preceq\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top};$
@@
$\leq\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}^{2}\mathbb{E}[(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))^{\top}]$
$\leq\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]-\gamma\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t}^{\prime})^{\top}+\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t})^{\top}]+\gamma^{2}\mathbb{E}[\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t}^{\prime})^{\top}]$
$\leq 2\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]-\gamma\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t}^{\prime})^{\top}+\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t})^{\top}]$
@@
Notice that the inequality on the fourth line follows from the assumption that $\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}\leq 1$ and the inequality on the fifth line follows from the fact that since $s_{t}$ and $s_{t}^{\prime}$ and both drawn from the stationary distribution, $\mathbb{E}[\phi(s_{t}^{\prime})\phi(s_{t}^{\prime})^{\top}]=\mathbb{E}[\phi(s_{t})\phi(s_{t})^{\top}]$.

### v3

@@ 3.3 Online estimation of the covariance matrix
Compared with this bootstrap-based procedure, our methodology is advantageous in the following ways: (i) the upper bound in Theorem 3.3 vanishes, as a function of $T$, at the rate of $O(T^{-\frac{1}{3}})$, which is faster than the $O(T^{-\frac{1}{4}}\log T)$ rate from the bound (18). (ii) The leading term in the upper bound in Theorem 3.3 entails a smaller problem-related coefficient than that of (18). (iii) Perhaps most importantly, in order to estimate the bootstrap distribution $\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T}$ with an $O(T^{-\frac{1}{4}})$ error rate, it is necessary to store the sample trajectory $\mathcal{S}_{T}$ and generate $O(T^{\frac{1}{2}})$ perturbed trajectories. Consequently, the bootstrap procedure has a time complexity of order $O(T^{\frac{3}{2}}d^{2})$ and a space complexity of order $O(T)$. When $d\ll T$, this amounts to a much higher computational burden than our online Algorithm 1.
3.4 Statistical inference
Combining the results from Sections 3.2 and 3.3, we can now propose a general methodology for constructing convex confidence sets for $\bm{\theta}^{\star}$ with guaranteed finite sample coverage. Based on the TD estimator $\overline{\bm{\theta}}_{T}$ (9) and the variance estimator $\widehat{\bm{\Lambda}}_{T}$ (15), and given any tolerance level $\delta\in(0,1)$, it is enough to pick any convex set $\mathcal{C}_{\delta}$ that has probability at least $1-\delta$ under the (random!) Gaussian distribution $\mathcal{N}(\overline{\bm{\theta}}_{T},\frac{\widehat{\bm{\Lambda}}_{T}}{T})$. That is, ssattisfying, the condition
$\mathbb{P}_{\bm{\theta}\sim\mathcal{N}(\overline{\bm{\theta}}_{T},\frac{\widehat{\bm{\Lambda}}_{T}}{T})}\left(\bm{\theta}\in\mathcal{C}_{\delta}\bigg{|}\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T}\right)\geq 1-\delta.$ (19)
Natural choices of $\mathcal{C}_{\delta}$ are (i) hyperctangles, leading to simultaneous confidence intervals of the form
$\mathcal{C}_{\delta}=\prod_{j=1}^{d}[(\overline{\bm{\theta}}_{T})_{j}-R/\sqrt{T},(\overline{\bm{\theta}}_{T})_{j}+R/\sqrt{T}],$ (20)
where $R>0$ is the quantile of the $\ell_{\infty}$ norm of a centered Gaussian random vectors with covariance $(\widehat{\bm{\Lambda}}_{T}/{T})$, which can be calculated via simulations; and (ii) ellipsoids or balls, for example
$\mathcal{C}_{\delta}=\left\{\bm{\theta}:\sqrt{T}\|\bm{\theta}-\overline{\bm{\theta}}_{T}\|_{\widehat{\bm{\Lambda}}_{T}^{-1}}^{2}\leq R\right\},$ (21)
where $R$ is the $1-\delta$ quantile of the $\chi^{2}(d)$ distribution.
The next result, whose proof is in Appendix B.4, derives a coverage rate for $\mathcal{C}_{\delta}$, which converges to the nominal level of $1-\delta$ at a rate of order $O(T^{-1/3})$, as a function of $T$ alone.
@@ 3.4 Statistical inference
Consider TD with Polyak-Ruppert averaging (7) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, for all $t>0$ with $\alpha=2/3$ and $\eta_{0}<1/2$.For a given significance level $\delta\in(0,1)$, it holds that
@@
uniformly over all convex sets $\mathcal{C}_{\delta}\subset\mathbb{R}^{d}$ satisfying (19).
We highlight two important features of the above result. First, the coverage guarantee holds uniformly over any choice of the convex set $\mathcal{C}_{\delta}$ meeting condition (19), including the common choices like (20) and (21). Secondly, the size (i.e., volume) of the confidence sets vanishes at the rate of $O(T^{-1/2})$, keeping all the other variable constant, suggesting that, in this regard, there is little or no room for improvements, at least in terms of rates.
4 Simulations
@@
We illustrate our theoretical results with some numerical experiments. [^3] The code for our experiments can be found at https://www.dropbox.com/scl/fo/tbw256dlx4gzxzgb4t3hq/AOUaiLFj79iSnc29dFTPpvQ?rlkey=nsgwlcbfx0b73uebwv8xlovag&st=1axbxaa1&dl=0 In order to test our proposed framework for statistical inference, we consider a member of a family of MDPs that are designed to be difficult to distinguish among each other. The specifics of the MDP are the same as described in Li et al. (2024), and we refer readers to Appendix D for the construction details. With $10,000$ independent trials, we examine the high-probability convergence of the $\ell_{2}$ norm of TD estimation errors, the rate by which they converge to their asymptotic Gaussian distributions, and the precision of our estimator for the covariance matrix. Figure 2 reveals that the empirical distribution of TD estimation error converges to the asymptotic Gaussian distribution slowly when $\alpha$ is set to ${1}/{2}$ as suggested by Samsonov et al. (2024). Notice that, for computational simplicity, the difference bewteen the empirical distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and the asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ is measured by
$d(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})):=\max_{1\leq j\leq d}\sup_{x\in\mathbb{R}}\left|\mathbb{P}(\sqrt{T}(\overline{\bm{\Delta}}_{T})_{j}\leq x)-\mathbb{P}((\bm{\Lambda}^{\star\frac{1}{2}}\bm{z})_{j}\leq x)\right|$
instead of the convex distance we invoked in the theoretical results.
@@
Figure 2: Distance between empirical and asymptotic distributions of the TD estimation error.
@@ 5 Discussion
We develop high-probability convergence guarantees and Berry-Esseen bounds for TD learning with linear function approximation and $i.i.d.$ samples in a non-asymptotic manner. We also propose an efficient way to construct a covariance estimator and derive its consistency rates. These results lead to a practicable methodology for constructing convex confidence sets for the target parameter, with its size vanishing at a parametric rate $O(T^{-1/2})$. An important direction for future research is to investigate whether the order $O(T^{-1/3})$ for the convergence speed of coverage rate in Theorem 3.4 are tight. Another interesting extension would be to derive specialized Berry-Esseen bounds for the sub-class of hyper-rectangles, where one would expect to find a better dependence on the dimension $d$ of the feature space, possibly of poly-logarithmic, as opposed to polynomial, order.
We believe our analysis and results can be adapted or extended to cover other iterative algorithms, such as stochastic gradient descent and SA methods. Recent analyses (see, e.g., Sheshukova et al., 2025; Samsonov et al., 2024) have demonstrated Berry-Esseen bounds for the convex distance (using an estimated variance and assuming i.i.d. data), as well as valid bootstrap schemes.
Acknowledgments and Disclosure of Funding
W. Wu and A. Rinaldo acknowledge partial support from the NIH under Grant R01NS121913. Gen Li was supported in part by the Chinese University of Hong Kong Direct Grant for Research. Y. Wei was supported in part by the NSF grants CCF-2106778, CCF-2418156 and CAREER award DMS-2143215. The authors thank Yang Peng for pointing us to Theorem A.19.
Appendix A Preliminaries
@@
A.1 Useful algebraic facts
Supplementary notation.
For any complex number $z\in\mathbb{C}$, we use $\mathsf{Re}(z)$ to denote its real part. If a matrix $\bm{M}$ is positive (semi-)definite, i.e. $\lambda_{\min}(\bm{M})>(\geq)0$, we write $\bm{M}\succ(\succeq)\bm{0}$; when two matrices $\bm{X}$ and $\bm{Y}$ satisfy $\bm{X}-\bm{Y}\succ(\succeq)\bm{0}$, we write $\bm{X}\succ(\succeq)\bm{Y}$ or equivalently $\bm{Y}\prec(\preceq)\bm{X}$. For a vector $\bm{x}\in\mathbb{R}^{d}$, we use $\{x_{i}\}_{1\leq i\leq d}$ to denote its entries; for a matrix $\bm{X}\in\mathbb{R}^{m\times n}$, we use $\{X_{ij}\}_{1\leq i\leq m,1\leq j\leq n}$ to denote its entries. The vectorization of matrix $\bm{X}\in\mathbb{R}^{m\times n}$, denoted as $\textbf{vec}(\bm{X})$, is defined as a vector in $\mathbb{R}^{mn}$ with entries $[\textbf{vec}(\bm{X})]_{(i-1)m+j}=\bm{X}_{ij}$ for every $1\leq i\leq m$ and $1\leq j\leq n$. For any matrices $\bm{X}\in\mathbb{R}^{m\times n}$, and $\bm{Y}\in\mathbb{R}^{p\times q}$, their Kronecker product, denoted as $\bm{X}\otimes\bm{Y}$, is a matrix of the size of $mp\times nq$, and can be written in blocked form as
With subscripts, it can be represented as $(\bm{X}\otimes\bm{Y})_{(i-1)m+j,(k-1)p+\ell}=X_{ik}Y_{j\ell}$. For a matrix sequence $\{\bm{F}_{t}\}_{t\in\mathbb{N}}$, a number sequence $\{f_{t}\}_{t\in\mathbb{N}}$ and a fixed matrix $\bm{X}$, if there exists a constant $\widetilde{C}$ independent of $t$ and $\bm{X}$ such that
$\|\bm{F}_{t}\|\leq\widetilde{C}f_{t}\|\bm{X}\|,\quad\text{and}\quad\mathsf{Tr}(\bm{F}_{t})\leq\widetilde{C}f_{t}\mathsf{Tr}(\bm{X})$ (22)
hold simultaneously, we denote for simplicity $\bm{F}_{t}=O(f_{t})\bm{X}$.
The following lemma captures some basic properties in linear algebra regarding the relationship between Frobenius and operator norms of matrices.
@@ A.1 Useful algebraic facts
    For any sequence of matrices $\bm{X}_{1},\bm{X}_{2},\ldots,\bm{X}_{n}\in\mathbb{R}^{d\times d}$:
@@
For any matrix $\bm{X}\in\mathbb{R}^{d\times d}$ and a symmetric matrix $\bm{Y}\in\mathbb{S}^{d\times d}$, if $\bm{Y}\succeq\bm{0}$, then
$\mathsf{Tr}(\bm{XY})=\mathsf{Tr}(\bm{YX})\leq\|\bm{X}\|\mathsf{Tr}(\bm{Y}).$
Lemma A.2 directly imply that for any matrix sequence $\{\bm{F}_{t}\}_{t\in\mathbb{N}}$ on $\mathbb{R}^{d\times d}$ and a fixed matrix $\bm{X}\succeq\bm{0}$, it can be guaranteed that
$\bm{F}_{t}\bm{X}=O(\|\bm{F}_{t}\|)\bm{X},\quad\text{and}\quad\bm{X}\bm{F}_{t}=O(\|\bm{F}_{t}\|)\bm{X}.$ (25)
The following lemma, due to Chen et al. (2020)[Lemma 4.1], bounds the difference between the inverse matrices and would be useful for our analysis of the plug-in variance estimator.
Lemma A.3.
Let $\bm{X},\bm{Y}$ be two invertible matrices that satisfy $\|\bm{X}^{-1}(\bm{Y}-\bm{X})\|\leq\frac{1}{2}$. Then the difference between $\bm{Y}^{-1}$ and $\bm{X}^{-1}$ can be bounded by
@@
The following lemma reveals several critical algebraic features of the matrix $\bm{A}$.
@@
Let $\bm{A}_{t},\bm{A}$ be defined as in (8) and (4) respectively, and $\bm{\Sigma}$, $\lambda_{0}$ and $\lambda_{\Sigma}$ be defined as in (5).Then the following features hold:
@@
$\mathbb{E}[\bm{A}_{t}^{\top}\bm{A}_{t}]\preceq\bm{A}+\bm{A}^{\top};$ (28)
$\bm{A}^{\top}\bm{A}\preceq\lambda_{\Sigma}(\bm{A}+\bm{A}^{\top});$ (29)
$\left\|\bm{I}-\eta\bm{A}\right\|\leq 1-\frac{1-\gamma}{2}\lambda_{0}\eta,\quad\forall\eta\in\left(0,\frac{1}{2\lambda_{\Sigma}}\right);$ (30)
$\|\bm{A}^{-1}\|\leq\frac{1}{\lambda_{0}(1-\gamma)}.$ (31)
@@
$-\left(\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top}\right)\preceq\bm{\phi}(s)\bm{\phi}(s^{\prime})^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s)^{\top}\preceq\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top};$
@@
$\preceq\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}^{2}\mathbb{E}[(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))^{\top}]$
$\preceq\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]-\gamma\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t}^{\prime})^{\top}+\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t})^{\top}]+\gamma^{2}\mathbb{E}[\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t}^{\prime})^{\top}]$
$\preceq 2\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]-\gamma\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t}^{\prime})^{\top}+\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t})^{\top}]$
@@
Notice that the inequality on the fourth line follows from the assumption that $\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}\leq 1$ and the inequality on the fifth line follows from the fact that since $s_{t}$ and $s_{t}^{\prime}$ and both drawn from the stationary distribution, $\mathbb{E}[\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t}^{\prime})^{\top}]=\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]$.
Proof of Equation (29).
It suffices to show that for any $\bm{x}\in\mathbb{R}^{d}$,
$\bm{x}^{\top}\bm{A}^{\top}\bm{Ax}\leq\lambda_{\Sigma}\bm{x}^{\top}(\bm{A}+\bm{A}^{\top})\bm{x}.$
By definition, the left hand side of this inequality is featured by
$\bm{x}^{\top}\bm{A}^{\top}\bm{Ax}=\|\bm{Ax}\|_{2}^{2}=\left\|\mathbb{E}[\bm{\phi}(s)(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime}))^{\top}\bm{x}]\right\|_{2}^{2},$
where the expectation is taken over the distribution $s\sim\mu,s^{\prime}\sim P(\cdot\mid s)$. For simplicity, denote
$f(s)=\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}[(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime}))^{\top}\bm{x}];$
then the law of total expectations yields
$\bm{x}^{\top}\bm{A}^{\top}\bm{Ax}=\left\|\mathbb{E}[f(s)\bm{\phi}(s)]\right\|_{2}^{2}.$
In what follows, we aim to show that
$\left\|\mathbb{E}[f(s)\bm{\phi}(s)]\right\|_{2}^{2}\leq\lambda_{\Sigma}\mathbb{E}[f^{2}(s)].$ (32)
Observe that for any vector $\bm{y}\in\mathbb{R}^{d}$, it can always be guaranteed that
$\mathbb{E}[f(s)-\bm{\phi}^{\top}(s)\bm{y}]^{2}\geq 0.$
A direct expansion yields
00 0 $\leq\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\bm{y}^{\top}\mathbb{E}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]\bm{y}$
$=\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\bm{y}^{\top}\bm{\Sigma}\bm{y}$
$\leq\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\lambda_{\Sigma}\|\bm{y}\|_{2}^{2},\quad\forall\bm{y}\in\mathbb{R}^{d}.$
By taking
$y=\frac{\mathbb{E}[f(s)\bm{\phi}(s)]}{\lambda_{\Sigma}},$
we obtain
$0\leq\mathbb{E}[f^{2}(s)]-\frac{\left\|\mathbb{E}[f(s)\bm{\phi}(s)]\right\|_{2}^{2}}{\lambda_{\Sigma}},$
and (32) follows immediately. In other words, it can be guaranteed for every $\bm{x}\in\mathbb{R}^{d}$ that
$\bm{x}^{\top}\bm{A}^{\top}\bm{Ax}\leq\lambda_{\Sigma}\mathbb{E}[f^{2}(s)].$
and we now proceed to bounding $\mathbb{E}[f^{2}(s)]$. The Jensen’s inequality implies
$\mathbb{E}[f^{2}(s)]$ $=\mathbb{E}_{s\sim\mu}\left\{\bm{x}^{\top}\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})]\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})]^{\top}\bm{x}\right\}$
$\leq\mathbb{E}_{s\sim\mu}\left\{\bm{x}^{\top}\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})][\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})]^{\top}\bm{x}\right\}$
$=\bm{x}^{\top}\mathbb{E}\left\{[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})][\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})]^{\top}\right\}\bm{x}$
$\leq\bm{x}^{\top}(\bm{A}+\bm{A}^{\top})\bm{x},$
where the last inequality follows from the same logic as in the proof of Equation (28). Equation (29) follows immediately.